please help me with this porblem...
A passenger train travels 392 kilometres in the same time that it takes a freight train to travel 322 kilometres. If the passenger train travels 20 km/hr faster than the freight train, find the rate of each train.
thanks.

Answer by stanbon(75887)

please help me with this porblem...
A passenger train travels 392 kilometres in the same time that it takes a freight train to travel 322 kilometres. If the passenger train travels 20 km/hr faster than the freight train, find the rate of each train.
---------------
Passenger train DATA:
distance = 392 km ; rate = x+20 km/h ; time = d/r = 392/(x+2) hrs
----------------------------------------
Freight train DATA:
distance = 322 km ; rate = x km/h ; time = d/r = 322/x hrs
----------------------------------------
Equation:
time = time
392/(x+2) = 322/x
392x = 322*x + 322*2
70x = 644
x = 9.2 km/h (rate of the freight train)
x+20 = 29.2 km/h (rate of the passenger train)
============================
Cheers,
Stan H.